# Verification and Synthesis of Compatible Control Lyapunov and Control Barrier Functions

### v1

@@ IV-B2 Cover many candidate states
With a feasible initial guess $V^{(0)}$ and $b^{(0)}$, by induction, the CLF $V^{(i)}$ and CBF $b^{(i)}$ are also compatible in each iteration.
V Numerical Simulation
In this section, we evaluate our proposed exact verification method and compatible CLF/CBF synthesis approach. We demonstrate that our verification can differentiate compatible and incompatible CLF/CBF by solving convex optimization problems. We show that our proposed synthesis approach construct CLF/CBF candidates by enlarging the compatible region while ensuring the compatibility. All case studies are written in Python based on Drake [43] and MOSEK [44], and are conducted on a laptop with an Intel Core i5 6300HQ CPU and 8GB RAM. The detailed parameter settings can be found in our code.
V-A Verification
We evaluate the proposed verification framework using the 2-D first order integrator example mentioned in [32]. In order to show the framework’s capability of differentiating incompatible and compatible CLF-CBF, we first construct an incompatible CBF and CLF and show that they fail to pass the verification. We then verify a compatible CBF/CLF pair.
We evaluate our verification approach on the 2D first-order integrator $\dot{\mathbf{x}}=\mathbf{u}$, where $\mathbf{x}\in\mathbb{R}^{2}$, $\mathbf{u}\in\mathcal{U}\subset\mathbb{R}^{2}$. The unsafe region $\mathcal{X}_{u}$ is represented by $\mathcal{X}_{u}=\{\mathbf{x}\;|\;\mathbf{x}_{1}^{2}+(\mathbf{x}_{2}-3)^{2}-1.5^{2}\leq 0\}$. We also consider the linear input constraint $\mathcal{U}=\{\mathbf{u}\;|\;|\mathbf{u}_{1}|\leq 3,|\mathbf{u}_{2}|\leq 3\}$. We let parameters $\kappa_{V}=0.01$ and $\kappa_{b}=0.01$.
V-A1 Incompatible Example
In this example, we use the following given pair of incompatible CLF and CBF:
[…]
The incompatibility is checked by finding a state $\hat{\mathbf{x}}=[0,5]^{T}$ at which there does not exist $\mathbf{u}\in\mathcal{U}$ satisfying CLF and CBF constraints. We then give this pair of CLF-CBF to the verification SOS program. The degrees of the Lagrangian polynomials are set to be $2$. The convex optimization (8) reports infeasibility, which is consistent with the incompatible state $[0,5]^{T}$.
V-A2 Compatible Example
We use the idea of CLBF [3] and give the verification framework a CLBF pair:
[…]
In this example, the polynomial degrees for compatible Lagrangian polynomials are set to be $2$. The solver solves the verification SOS program successfully, which certifies the compatibility of the CLF and CBF in (V-A2) are compatible, and the given CBF is correct.
The results for these two examples show that our proposed verification approach can verify compatibility of CBF-CLF, while also differentiating between compatible and incompatible CBF-CLF pairs.
V-B Synthesis
We next evaluate our proposed bilinear alternation synthesis method. We investigate a linear example, a nonlinear example [45] and a power converter example [33]. We set parameters $\kappa_{V}=0.01$, and $\kappa_{b}=0.01$. Since the compatible region may infinitely enlarged if $\mathcal{X}_{s}$ is an open set, then the synthesis loop may not terminate. Hence, we set maximum number of iterations as a hard stopping condition (MaxIter=25) during experiments.
V-B1 Linear Example
For the linear toy example, we use the following system dynamics:
$\begin{bmatrix}\dot{\mathbf{x}}_{1}\\ \dot{\mathbf{x}}_{2}\end{bmatrix}=\begin{bmatrix}1&2\\ -2&3\end{bmatrix}\begin{bmatrix}\mathbf{x}_{1}\\ \mathbf{x}_{2}\end{bmatrix}+\begin{bmatrix}1&0\\ 0&1\end{bmatrix}\begin{bmatrix}\mathbf{u}_{1}\\ \mathbf{u}_{2}\end{bmatrix}.$ (19)
The obstacle is represented as $\mathcal{X}_{u}=\{\mathbf{x}\;|\;\mathbf{x}_{1}+\mathbf{x}_{2}+3\leq 0\}$ We construct a compatible CBF and CLF using a heuristic method that solves a linear quadratic regulator (LQR) problem in a neighborhood of an equilibrium point to initialize CLF/CBF. We set the cost matrices of LQR as $Q=I_{2},R=I_{2}$, where $I_{2}$ is $2\times 2$ identity matrix.
Figure 1: This figure presets the compatible CLF-CBF synthesis results for all the investigated examples. In these plots, unsafe regions (UR) are marked as red, initial compatible regions (ICR) are marked as gray, initial compatible boundaries (ICB) are marked as dashed lines, synthesized compatible regions (SCB) are marked as green, and synthesized compatible boundaries (SCB) are marked as solid lines. The compatible regions are enlarged for all the investigated examples. Figure 1(a) and 1(b) show the results for linear example and nonlinear example, respectively. In these two examples, the synthesis loop stops after maximum number of iterations (25 iterations). Figure 1(c) shows the synthesis result for power converter example. The synthesis stops after 7 iterations because the compatible region can not be enlarged further according to optimization (16)
Figure 1(a) shows the initial compatible region and the compatible region after synthesis. We observe that our alternating synthesis method can enlarge the compatible region and incorporates the initial compatible region. The synthesis loop stops after maximum number of iterations.
V-B2 Nonlinear Example
In this example, the following dynamics [21] is used:
$\begin{bmatrix}\dot{\mathbf{x}}_{1}\\ \dot{\mathbf{x}}_{2}\end{bmatrix}=\begin{bmatrix}0\\ -\mathbf{x}_{1}+\frac{\mathbf{x}_{1}^{3}}{6}\end{bmatrix}+\begin{bmatrix}1\\ -1\end{bmatrix}\begin{bmatrix}u\end{bmatrix}.$ (20)
We pick $\mathbf{x}_{eq}=[0,0]^{T},u_{eq}=0$ as the equilibrium point and linearize the system around this point to compute the LQR. The obstacle for this example is the same as linear toy example. The cost matrices of LQR are set as $Q=I_{2},R=1$.The synthesis results for this example are shown by Figure 1(b). We observe that the proposed alternating synthesis can enlarge the compatible region in this example.
V-B3 Power Converter Example
The dynamical model of the power converter is given in [33] as follows
$\begin{bmatrix}\dot{\mathbf{x}}_{1}\\ \dot{\mathbf{x}}_{2}\\ \dot{\mathbf{x}}_{3}\end{bmatrix}=\begin{bmatrix}-0.05\mathbf{x}_{1}-57.9\mathbf{x}_{2}+0.00919\mathbf{x}_{3}\\ 1710\mathbf{x}_{1}+314\mathbf{x}_{3}\\ -0.271\mathbf{x}_{1}-314\mathbf{x}_{2}\end{bmatrix}+\\ \begin{bmatrix}0.05-57.9\mathbf{x}_{2}&-57.9\mathbf{x}_{3}\\ 1710+1710\mathbf{x}_{1}&0\\ 0&1710+1710\mathbf{x}_{1}\end{bmatrix}\begin{bmatrix}\mathbf{u}_{1}\\ \mathbf{u}_{2}\end{bmatrix}.$ (21)
The unsafe region is presented as:
$\mathcal{X}_{u}=\{\mathbf{x}|$ $-\mathbf{x}_{1}+0.2\leq 0,\mathbf{x}_{1}+0.8\leq 0,$
$-(\mathbf{x}_{2}-0.001)^{2}-\mathbf{x}_{3}^{2}+1.2^{2}\leq 0\}.$ (22)
The equilibrium point of the system is $\mathbf{x}_{eq}=[0,0,0]^{T},\mathbf{u}_{eq}=[0,0]^{T}$. We linearize the system near the equilibrium point and compute the LQR. We set the cost matrices of LQR as $Q=I_{3},R=I_{2}$, where $I_{3}$ is $3\times 3$ identity matrix. The initialized and synthesised compatible regions for power converter example are shown by Figure 1(c). The proposed synthesis method enlarges the compatible region for the power converter. The syntehsis loop stops after 7 number of iterations because the compatible region can not be enlarged further according to optimization problem (16). The returned CLF-CBF are verified to be compatible. Together with Figure 1(a) and Figure 1(b), we can conclude that the proposed alternating synthesis method enlarges the compatible region and returns compaitble CLF-CBF for all the investigated examples.
@@ VI Conclusion
In this paper, for control affine systems both with and without input limits, we studied the verification and synthesis of compatible CBF and CLF independent from any nominal controllers. We presented sufficient and necessary conditions for compatible CBF and CLF via Farkas’ Lemma and algebraic-geometric Positivstellensatz, and demonstrated that these conditions can be checked by solving an SOS program. In order to reduce the number of SOS polynomials in the SOS program, we further simplified the necessary and sufficient condition using S-procedure and formulated a more tractable SOS program for compatibility verification. We next proposed an alternating method to synthesize compatible CLF and CBFs based on the verification framework. In each iteration, the algorithm first takes the CLF and CBF from the previous iteration and computes the Lagrangian polynomials by solving the verification SOS program. Then, it takes the Lagrangian polynomials and updates the CLF and CBF to obtain a larger compatible region. Finally, we evaluated our proposed approach with a linear system, a nonlinear system and nonlinear power converter. In our future work, we will extend the verification and synthesis framework to high-relative degree and non-convex scenarios.

### v2

@@ IV-B Compatible CBF and CLF synthesis
With a feasible initial guess $V^{(0)}$ and $h^{(0)}$, by induction, the CLF $V^{(i)}$ and CBF $h^{(i)}$ are also compatible in each iteration.
V Results
In this section, we apply our algorithm to different dynamical systems. We choose a toy system whose results can be easily visualized on 2D plots, and a 13-state quadrotor robot to showcase the scalability of our approach. All optimization problems are solved via Mosek solver [47]. We report the CLF/CBF and Lagrangian degrees, together with the computation time in Table I.
Polynomial degree Computation time
$V(\mathbf{x})$ $h(\mathbf{x})$ $s_{0}(\mathbf{x},\mathbf{y})$ $s_{1}(\mathbf{x},\mathbf{y})$ $s_{2}(\mathbf{x},\mathbf{y})$ $s_{3}(\mathbf{x},\mathbf{y})$ $p(\mathbf{x})$ $q(\mathbf{x})$ Problem (8) Problem (15)
Toy 2 2 2 2 4 4 0 0 0.05s 0.08s
Quadrotor 2 2 1 1 4 4 0 0 213s 149s
TABLE I: Degrees of CLF/CBF, Lagrangian polynomials and computation time.
V-A Toy system
We consider the following toy system (from [17])
$\dot{\theta}=$ $u$ (16a)
$\dot{\gamma}=$ $-\sin\theta-u$ (16b)
$-1\leq$ $u\leq 1,$ (16c)
where the unsafe region is
$\mathcal{X}_{u}=\{(\theta,\gamma)|\sin\theta+\cos\theta+\gamma+1\leq 0\}.$ (17)
First, we convert the system dynamics to the polynomial form, by introducing new variables $s=\sin\theta,c=\cos\theta$, with the state defined as $\mathbf{x}=[x_{1},x_{2},x_{3}]=[s,c-1,\gamma]$. Note that we subtract $1$ from $\cos\theta$ to make sure that the equilibrium $\theta=0,\gamma=0$ is at $\mathbf{x}=\mathbf{0}$. The polynomial dynamics is
$\dot{\mathbf{x}}=\begin{bmatrix}0\\ 0\\ -x_{1}\end{bmatrix}+\begin{bmatrix}x_{2}+1\\ -x_{1}\\ -1\end{bmatrix}u$ (18a)
$-1\leq u\leq 1,$ (18b)
with the constraint $\sin^{2}\theta+\cos^{2}\theta=1$ as the following algebraic constraint on the state $\mathbf{x}$
$e(\mathbf{x})=x_{1}^{2}+(x_{2}+1)^{2}-1=0.$ (18c)
The unsafe region is $\mathcal{X}_{u}=\{\mathbf{x}|x_{1}+x_{2}+x_{3}+2\leq 0\}$.
We synthesize compatible CLF/CBF through algorithm 1. We choose the initial CLF as $V_{\text{initial}}(\mathbf{x})=10(x_{1}^{2}+x_{2}^{2}+x_{3}^{2})$ and the initial CBF as $h_{\text{initial}}(\mathbf{x})=1-V_{\text{initial}}(\mathbf{x})$. We plot the result in Fig.2(a) (note that the plot is in the original $(\theta,\gamma)$ space). We also draw the incompatible region, namely the region where the CLF constraint $L_{f}V(\mathbf{x})+L_{g}V(\mathbf{x})u+\kappa_{V}V(\mathbf{x})\leq 0$ and the CBF constraint $L_{f}h(\mathbf{x})+L_{g}h(\mathbf{x})u+\kappa_{h}h(\mathbf{x})\geq 0$ don’t share a common solution for $u$ within the input limit $u\in[-1,1]$. Fig 2(a) shows that
  1. 1.
    Algorithm 1 can successfully find compatible CLF and CBF that expand the compatible region $\{\mathbf{x}|V(\mathbf{x})\leq 1,h(\mathbf{x})\geq 0\}$, which grows from the small ellipsoid in the center to the large green region.
  2. 2.
    The synthesized compatible region is very tight, as the green compatible region almost touches the grey obstacle region and the cyan incompatible region. This showcases that our S-procedure formulation (8) can generate very tight solution with low-degree Lagrangian multipliers (the Lagrangians are of degree no larger than 4).
(a) Synthesize compatible CLF/CBF. The grey region is the obstacle. The cyan region is where CLF is incompatible with the CBF. The compatible region grows from the small circle $\{\mathbf{x}|V_{\text{initial}}(\mathbf{x})\leq 1,h_{\text{initial}}(\mathbf{x})\geq 0\}$ to the large green region. The green compatible region doesn’t touch the grey obstacle or the cyan incompatible region.
(b) Synthesize CLF and CBF separately using approaches from previous works. They are incompatible with each other, as demonstrated in the highlighted erroneous region, where the set $\{\mathbf{x}|V(\mathbf{x})\leq 1,h(\mathbf{x})\geq 0\}$ intersects with the cyan incompatible region in which there doesn’t exist a control action to satisfy CLF constraint and CBF constraint simultaneously.
Figure 2: Results of jointly synthesizing compatible CLF/CBF (Fig 2(a), our approach), and separately synthesizing incompatible CLF/CBF (Fig 2(b), previous approach) for the nonlinear toy example.
As a comparison, we use the previous approach [22] to synthesize the CLF and CBF separately (The CBF is set as $h(\mathbf{x})=x_{1}+x_{2}+x_{3}+2$, such that $\{\mathbf{x}\;|\;h(\mathbf{x})\geq 0\}$ is exactly the safe region $\mathcal{X}_{s}$). As shown in Fig 2(b), ignoring the compatibility condition during CLF/CBF synthesis leads to the existence of the erroneous region, in which the CLF constraint and CBF constraint cannot be satisfied simultaneously, causing the CLF-CBF-QP controller to fail at the runtime.
V-B Quadrotor
To showcase the scalability of our approach, we apply it to a 3D quadrotor with 13 states, including quaternion for its body orientation, the position, linear velocity and angular velocity. The input $\mathbf{u}\in\mathbb{R}^{4}$ is the rotor thrust. We have the extra constraint $\mathbf{q}^{T}\mathbf{q}=1$ on the unit-length quaternion $\mathbf{q}$. The polynomial dynamics of the quadrotor is described in [48]. The obstacle is the ground at height $-0.5m$.
We synthesize the compatible CLF/CBF through Algorithm 1333For better numerics, we use SDSOS constraint instead of SOS constraint on the Lagrangian multiplier $s_{2}(\mathbf{x},\mathbf{y}),s_{3}(\mathbf{x},\mathbf{y})$, where SDSOS is a stronger convex constraint than SOS [49]., with the initial CLF computed from the LQR solution, and the initial CBF as $1-V_{\text{initial}}(\mathbf{x})$. Note that the CBF depends on both the position and velocity, hence the relative-degree of the CBF equals to 1. After optimizing the CLF/CBF, we simulate the quadrotor with the CLF-CBF-QP controller, whose cost is to minimize the thrust $\mathbf{u}^{T}\mathbf{u}$. We start the simulation from different initial states within the compatible region, as shown in Fig 1. Since the QP objective is to minimize the thrust, the quadrotor first descends (with small thrusts), and then the CBF constraint will prevent it from hitting the ground and the quadrotor lifts up to the goal. We also draw the CLF/CBF trajectories along the simulation in Fig 3. The CLF values converge to 0 while the CBF values stay above 0.
Figure 3: CLF/CBF trajectories of the quadrotor while simulated with the CLF-CBF-QP controller.
@@ VI Conclusion
In this paper, for control affine systems both with and without input limits, we studied the verification and synthesis of compatible CBF and CLF independent from any nominal controllers. We presented necessary and sufficient conditions for compatible CBF and CLF via Farkas’ Lemma and algebraic-geometric Positivstellensatz, and demonstrated that these conditions can be verified by solving a SOS program. In order to reduce the number of sos polynomials, we further simplified the necessary and sufficient condition using S-procedure and formulated a more tractable SOS program for compatibility verification. We solve a sequence of SOS programs so as to synthesize CLF/CBF with a large compatible region. Finally, we evaluated our proposed approach on a toy system and a quadrotor.
In our future work, it is natural to extend the verification and synthesis framework to high relative degree CBF [50]. Moreover, in this work we used a single polynomial function as the CBF. It would require a very high degree polynomial to capture the safe region with a complicated shape, resulting in a large-size optimization problem and poor numerics. To remedy this, we will consider using the minimal of several compatible low-degree polynomials as the CBF, which we intend to address in our follow-up work.